OpenStudy (anonymous):

If y varies inversely as x, and y = 4 as x = 8, find y for the x-value of 2.

OpenStudy (anonymous):

A. 16 B. 8 C. 14 D. 10

OpenStudy (lynfran):

y=k/x

OpenStudy (anonymous):

yes, i know. but i don't know how to use it here.

OpenStudy (anonymous):

i know that k=4, but that's that.

OpenStudy (lynfran):

ok so 4=k/8....multiply both side of the equation by 8 we get 32=k

OpenStudy (anonymous):

okay...

OpenStudy (lynfran):

when x=2, y=32/2

OpenStudy (anonymous):

32/2=16.

OpenStudy (anonymous):

thanks!!!

OpenStudy (lynfran):

correct

OpenStudy (anonymous):

thanks so much. you really helped me. alot.

OpenStudy (lynfran):

ur welcome